A physician orders 1.5 g IM injection of an antibiotic. The dose needs to be administered in a 2 mL volume. The package insert for a 6 g vial states " add 8.6 mL of diluent for a final concentration of 1g/2mL. How many mLs of diluent should be added to make the requested concentration.
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

       Formula

           Volume(ml) = Desired dose / Dose in hand * Quantity.

   Here

       Desired dose (D) = 1.5 gm

       Dose in hand (H) = 1 gm

           Quantity (Q) = 2 ml

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 1.5/1 * 2

                                                      = 1.5 * 2

                                                      = 3 ml